# Local Dimension Spectrum for Dominated Irreducible Planar Self-Affine Sets

### v1

@@ 3 Differentiability of the Pressure
Let $(q_{k},s_{k})_{k\in\mathbb{N}}$ be a sequence of points in ×$\mathbb{R}\times(0,\infty)\setminus\{1,2\}$ converging to $(q,s)\in\mathbb{R}\times(0,\infty)\setminus\{1,2\}$. Then $\psi^{q_{k},s_{k}}(\mathtt{i})^{1/|\mathtt{i}|}\to\psi^{q,s}(\mathtt{i})^{1/|\mathtt{i}|}$ uniformly in $\Sigma_{*}$ as $k\to\infty$.
@@
Following notation of the proof of Lemma 3.1, the estimates (3.1) and (3.2) give
@@
for all $\mathtt{i}\in\Sigma_{*}$ and $k\in\mathbb{N}$. The claim follows. ∎
Before proving the differentiability, let us recall some further facts from convex analysis. Let $U\subset\mathbb{R}$ be an open set and let $f\colon U\to\mathbb{R}$ be convex. Let $D\subset U$ be the set of points where $f$ is differentiable. If $z_{1},z_{2}\in D$ and $x\in U$ such that $z_{1}\leqslant x\leqslant z_{2}$, then, by [20, Theorem 24.1], $f^{\prime}(z_{1})\leqslant G\leqslant f^{\prime}(z_{2})$ for all sub-derivatives $G$ at $x$. It also follows that the set $U\setminus D$ is at most countable and $f^{\prime}$ is continous on $D$; see [20, Theorem 25.3]. In particular, $D$ is dense in $U$ and if $f$ is differentiable on $U$, then it is continuously differentiable on $U$.
@@
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be dominated tuple of contractive matrices and $\mu\in\mathcal{M}(\Sigma)$ be a quasi-Bernoulli measure. Then the function $(q,s)\mapsto P(\psi^{q,s})$ is differentiable on ×$\mathbb{R}\times(0,\infty)\setminus\{1,2\}$.
@@
We prove the result only on ×$\mathbb{R}\times(0,2)\setminus\{1\}$; the case ×$\mathbb{R}\times(2,\infty)$ is left to the reader. To simplify notation, we write $Q(q,s)=P(\psi^{q,s})$ for all $(q,s)\in\mathbb{R}\times(0,2)\setminus\{1\}$. To see that $Q$ is differentiable on ×$\mathbb{R}\times(0,2)\setminus\{1\}$, it suffices to show that both partial derivatives of $Q$ exist at each point of ×$\mathbb{R}\times(0,2)\setminus\{1\}$. Indeed, assuming this is the case, then using Proposition 3.3 combined with Lemmata 2.4, 2.3, and 3.4 it is straightforward to prove the continuity of the partial derivatives which then implies the differentiability.
@@
Let $\nu_{k}^{-}$ and $\nu_{k}^{+}$ be the equilibrium states associated to $\psi^{q_{k}^{-},s_{0}}$ and $\psi^{q_{k}^{+},s_{0}}$, respectively. Then, by Lemmata 3.4 and 2.4, any limit point of the sequences $(\nu_{k}^{-})_{k\in\mathbb{N}}$ and $(\nu_{k}^{+})_{k\in\mathbb{N}}$ must be the unique equilibrium state $\nu$ of $\psi^{q_{0},s_{0}}$. Thus, $\nu_{k}^{-}\to\nu$ and $\nu_{k}^{+}\to\nu$ by the compactness of $\mathcal{M}_{\sigma}(\Sigma)$. Let $G_{1}$ be a sub-derivative of $q\mapsto Q(q,s_{0})$ at $q_{0}$. It follows from Proposition 3.3 that
@@
Now we show that other partial derivative exists. By Lemma 3.1(6), we know that the partial derivative $\partial_{s}Q(q_{0},s)$ exists for all, except possibly at countably many points on $[0,\infty)$. Relying on this, choose two sequences $(s_{k}^{-})_{k\in\mathbb{N}}$ and $(s_{k}^{+})_{k\in\mathbb{N}}$ of points in $(0,2)\setminus\{1\}$ with $\lceil s_{k}^{-}\rceil=\lceil s_{k}^{+}\rceil=\lceil s_{0}\rceil$, $s_{k}^{-}\uparrow s_{0}$, and $s_{k}^{+}\downarrow s_{0}$ as $k\to\infty$ so that the partial derivatives $\partial_{s}Q(q_{0},s_{k}^{-})|_{s=s_{k}^{-}}$ and $\partial_{s}Q(q_{0},s_{k}^{+})|_{s=s_{k}^{+}}$ exist.
@@ 4 Symbolic Spectrum
For each $q\neq 1$ let $s_{q}\in[0,\infty)$ be as in Lemma 3.1(4). The $L^{q}$-spectrum is the function $\tau\colon\mathbb{R}\to\mathbb{R}$ defined by $\tau(q)=(q-1)s_{q}$ for all $q\in\mathbb{R}\setminus\{1\}$ and $\tau(1)=0$. The symbolic multifractal spectrum is the function $f\colon[0,\infty)\to\mathbb{R}$ defined by
@@
We say that $\tau$ and $f$ form a Legendre transform pair at $(q,s)$ if $f(s)=qs-\tau(q)$. In this section, we show that $\tau$ and $f$ form a Legendre transform pair at $(q,\tau^{\prime}(q))$. To that end, let us begin by verifying the existence of $\tau^{\prime}$ and reveal its connection to the Lyapunov cross-dimension.
@@
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be dominated tuple of contractive matrices, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, and $q\neq 1$. Then the $L^{q}$-spectrum $\tau$ is continuously differentiable on a neighborhood of $q$ with
@@
Fix $q_{0}\neq 1$ and, by Lemma 3.1(4), let $s_{q_{0}}\geqslant 0$ be such that $P(\psi^{q_{0},s_{q_{0}}})=0$. Note that, by the definition of $\tau$, $s_{q_{0}}=\tau(q_{0})/(q_{0}-1)$. Recalling Propositions 3.5 and 3.3, implicit differentiation gives
@@
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be dominated tuple of contractive matrices, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, and $q\neq 1$. If $\tau(q)/(q-1)$ and $q\tau^{\prime}(q)-\tau(q)$ are contained in the same interval, $(0,1)$, $(1,2)$ or $(2,\infty)$, then
@@
$-\frac{h(\nu)}{\lambda_{1}(\nu)}=-q\frac{h(\mu,\nu)}{\lambda_{1}(\nu)}-\tau(q)=q\tau^{\prime}(q)-\tau(q).$
@@
$-\frac{2h(\nu)}{\lambda_{1}(\nu)+\lambda_{2}(\nu)}=-q\frac{2h(\mu,\nu)}{\lambda_{1}(\nu)+\lambda_{2}(\nu)}-\tau^{\prime}(q)=q\tau^{\prime}(q)-\tau(q).$
@@
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be dominated tuple of contractive matrices, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, and $q\neq 1$. If $\tau^{\prime}(q)$, $\tau(q)/(q-1)$, and $q\tau^{\prime}(q)-\tau(q)$ are contained in the same interval, $(0,1)$, $(1,2)$ or $(2,\infty)$, then $\tau^{\prime}(q)\leqslant\tau(q)/(q-1)$ and
@@ 5 Exact Dimensionality
for all $\mathtt{i}=i_{1}i_{2}\cdots\in\Sigma$. Here $\varphi_{\mathtt{i}}=\varphi_{i_{1}}\circ\cdots\circ\varphi_{i_{n}}$ for all $\mathtt{i}=i_{1}\cdots i_{n}\in\Sigma_{n}$ and $n\in\mathbb{N}$. It is easy to see that $\pi(\Sigma)=X$. We say that $X$ satisfies the strong separation condition if $\varphi_{i}(X)\cap\varphi_{j}(X)=\emptyset$ whenever $i\neq j$. As $\pi([\mathtt{i}])=\varphi_{\mathtt{i}}(X)$ for all $\mathtt{i}\in\Sigma_{*}$, the strong separation condition is characterized by the property that the canonical projection $\pi$ is one-to-one. We use the convention that whenever we speak about a self-affine set, then it is automatically accompanied with a tuple of affine maps which defines it. This makes it possible to write that e.g. “a self-affine set $X$ is dominated” which obviously then means that “the associated tuple $\mathsf{A}$ of matrices is dominated”.
Let $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ be a quasi-Bernoulli measure. If $X$ is dominated, then, by [5, Theorem 2.6] and [21, Theorem 2.2],
$\operatorname{dim_{loc}}(\pi_{*}\nu,\pi(\mathtt{i}))=\operatorname{\underline{dim}_{H}}(\pi_{*}\nu)\leqslant\operatorname{dim_{L}}(\nu)$
for $\nu$-almost all $\mathtt{i}\in\Sigma$. In other words, canonical projections of quasi-Bernoulli measures, which we call quasi-Bernoulli measures on $X$, are exact-dimensional. We say that $\mathsf{A}=(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ is irreducible if there is no line in $\mathbb{R}^{2}$ which is invariant under all of the matrices in $\mathsf{A}$. If $\mathsf{A}$ is dominated such that there is a finite set of lines in $\mathbb{R}^{2}$ invariant under all of the matrices in $\mathsf{A}$, then, by [7, Lemma 2.10], $\mathsf{A}$ is not irreducible. Therefore, if $X$ is dominated and irreducible, and satisfies the strong separation condition, then, by [3, Theorem 1.2 and the associated footnote],
@@
for all quasi-Bernoulli measures $\nu\in\mathcal{M}_{\sigma}(\Sigma)$.
In this section, by generalizing the approach used in [14, §4.1], we study the local dimension of a quasi-Bernoulli measure on a self-affine set with respect to another measure. We first recall the following lemma.
@@
The proof is the same as that of [14, Lemma 4.2] where the result is formulated in the case of positive matrices. ∎
@@
Let $X$ be a dominated planar self-affine set satisfying the strong separation condition. If $\mu\in\mathcal{M}(\Sigma)$ is fully supported quasi-Bernoulli and $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ is ergodic, and if
@@
satisfies $\nu\times\nu_{F}(\Gamma_{\kappa})>1-\kappa$. As $\nu\times\nu_{F}$ is $T$-invariant and ergodic, Birkhoff’s ergodic theorem implies that
@@
$=\nu\times\nu_{F}(\Gamma_{\kappa})>1-\kappa$
for $\nu\times\nu_{F}$-almost all $(\mathtt{i},V)\in\Sigma\times\mathbb{RP}^{1}$. Write $r_{n}(\mathtt{i})=\alpha_{2}(\mathtt{i}|_{n})/\alpha_{1}(\mathtt{i}|_{n})$ for all $\mathtt{i}\in\Sigma$ and $n\in\mathbb{N}$, and note that, by domination, there exists $n_{0}\in\mathbb{N}$ such that $r_{n}(\mathtt{i})<r_{0}$ for all $\mathtt{i}\in\Sigma$ and $n\geqslant n_{0}$. By letting $\kappa\downarrow 0$, it follows from the above estimate that
@@
Let $\varrho_{2}>0$ be as in Lemma 5.1 and notice that, as $\alpha_{2}$ is almost-multiplicative by domination, there exists a constant $C\geqslant 1$ such that
11 11 $\geqslant\frac{\log\varrho_{2}\alpha_{2}(\mathtt{i}|_{n_{k+1}})}{\log\varrho_{2}\alpha_{2}(\mathtt{i}|_{n_{k}})}\geqslant\frac{\log\varrho_{2}\alpha_{2}(\mathtt{i}|_{n_{k}})+\log\alpha_{2}(\sigma^{n_{k}}\mathtt{i}|_{n_{k+1}-n_{k}})+\log C}{\log\varrho_{2}\alpha_{2}(\mathtt{i}|_{n_{k}})}$

### v2

@@ 3 Differentiability of the Pressure
Let $(q_{k},s_{k})_{k\in\mathbb{N}}$ be a sequence of points in $\mathbb{R}\times(0,\infty)\setminus\{1,2\}$ converging to $(q,s)\in\mathbb{R}\times(0,\infty)\setminus\{1,2\}$. Then $\psi^{q_{k},s_{k}}(\mathtt{i})^{1/|\mathtt{i}|}\to\psi^{q,s}(\mathtt{i})^{1/|\mathtt{i}|}$ uniformly in $\Sigma_{*}$ as $k\to\infty$.
@@
Using the notation from the proof of Lemma 3.1, the estimates (3.1) and (3.2) imply
@@
for all $\mathtt{i}\in\Sigma_{*}$ and $k\in\mathbb{N}$. This proves the claim. ∎
Before proving the differentiability, let us recall some further facts from convex analysis. Let $U\subset\mathbb{R}$ be an open set and let $f\colon U\to\mathbb{R}$ be convex. Let $D\subset U$ be the set of points where $f$ is differentiable. If $z_{1},z_{2}\in D$ and $x\in U$ such that $z_{1}\leqslant x\leqslant z_{2}$, then, by [33, Theorem 24.1], $f^{\prime}(z_{1})\leqslant G\leqslant f^{\prime}(z_{2})$ for all sub-derivatives $G$ at $x$. It also follows that the set $U\setminus D$ is at most countable and $f^{\prime}$ is continuous on $D$; see [33, Theorem 25.3]. In particular, $D$ is dense in $U$ and if $f$ is differentiable on $U$, then it is continuously differentiable on $U$.
@@
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be a dominated tuple of contractive matrices and $\mu\in\mathcal{M}(\Sigma)$ be a quasi-Bernoulli measure. Then the function $(q,s)\mapsto P(\psi^{q,s})$ is differentiable on $\mathbb{R}\times(0,\infty)\setminus\{1,2\}$.
@@
We prove the result only on $\mathbb{R}\times(0,2)\setminus\{1\}$; the case $\mathbb{R}\times(2,\infty)$ is left to the reader. To simplify notation, we write $Q(q,s)=P(\psi^{q,s})$ for all $(q,s)\in\mathbb{R}\times(0,2)\setminus\{1\}$. To see that $Q$ is differentiable on $\mathbb{R}\times(0,2)\setminus\{1\}$, it suffices to show that both partial derivatives of $Q$ exist at each point of $\mathbb{R}\times(0,2)\setminus\{1\}$. Indeed, assuming this is the case, then using Proposition 3.2 combined with Lemmata 2.4, 2.3, and 3.3 it is straightforward to prove the continuity of the partial derivatives which then implies the differentiability.
@@
Let $\nu_{k}^{-}$ and $\nu_{k}^{+}$ be the equilibrium states associated to $\psi^{q_{k}^{-},s_{0}}$ and $\psi^{q_{k}^{+},s_{0}}$, respectively. Then, by Lemmata 3.3 and 2.4, any limit point of the sequences $(\nu_{k}^{-})_{k\in\mathbb{N}}$ and $(\nu_{k}^{+})_{k\in\mathbb{N}}$ must be the unique equilibrium state $\nu$ of $\psi^{q_{0},s_{0}}$. Thus, $\nu_{k}^{-}\to\nu$ and $\nu_{k}^{+}\to\nu$ by the compactness of $\mathcal{M}_{\sigma}(\Sigma)$. Let $G_{1}$ be a sub-derivative of $q\mapsto Q(q,s_{0})$ at $q_{0}$. It follows from Proposition 3.2 that
@@
Now we show that the other partial derivative exists. By Lemma 3.1(6), we know that the partial derivative $\partial_{s}Q(q_{0},s)$ exists for all, except possibly at countably many points on $0,\infty)$. Relying on this, choose two sequences $(s_{k}^{-})_{k\in\mathbb{N}}$ and $(s_{k}^{+})_{k\in\mathbb{N}}$ of points in $(0,2)\setminus\{1\}$ with $\lceil s_{k}^{-}\rceil=\lceil s_{k}^{+}\rceil=\lceil s_{0}\rceil$, $s_{k}^{-}\uparrow s_{0}$, and $s_{k}^{+}\downarrow s_{0}$ as $k\to\infty$ so that the partial derivatives $\partial_{s}Q(q_{0},s_{k}^{-})|_{s=s_{k}^{-}}$ and $\partial_{s}Q(q_{0},s_{k}^{+})|_{s=s_{k}^{+}}$ exist.
@@ 4 Symbolic Spectrum
For each $q\neq 1$ let $s_{q}\in[0,\infty)$ be as in Lemma 3.1(4). The symbolic $L^{q}$-spectrum is the function $\tau\colon\mathbb{R}\to\mathbb{R}$ defined by $\tau(q)=(q-1)s_{q}$ for all $q\in\mathbb{R}\setminus\{1\}$ and $\tau(1)=0$. The symbolic multifractal spectrum is the function $f\colon[0,\infty)\to\mathbb{R}$ defined by
@@
We say that $\tau$ and $f$ form a Legendre transform pair at $(q,s)$ if $f(s)=qs-\tau(q)$. In this section, we show that $\tau$ and $f$ form a Legendre transform pair at $(q,\tau^{\prime}(q))$. To that end, following [10, §3], let us begin by verifying the existence of $\tau^{\prime}$ and reveal its connection to the Lyapunov cross-dimension.
@@
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be a dominated tuple of contractive matrices, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, and $q\neq 1$. Then the symbolic $L^{q}$-spectrum $\tau$ is continuously differentiable on a neighborhood of $q$ with
@@
Fix $q_{0}\neq 1$ and, by Lemma 3.1(4), let $s_{q_{0}}\geqslant 0$ be such that $P(\psi^{q_{0},s_{q_{0}}})=0$. Note that, by the definition of $\tau$, $s_{q_{0}}=\tau(q_{0})/(q_{0}-1)$. Recalling Propositions 3.4 and 3.2, implicit differentiation gives
@@
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be a dominated tuple of contractive matrices, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, and $q\neq 1$. If $\tau(q)/(q-1)$ and $q\tau^{\prime}(q)-\tau(q)$ are contained in the same interval, $(0,1)$, $(1,2)$ or $(2,\infty)$, then
@@
$\frac{h(\nu)}{\lambda_{1}(\nu)}=-q\frac{h(\mu,\nu)}{\lambda_{1}(\nu)}-\tau(q)=q\tau^{\prime}(q)-\tau(q).$
@@
$\frac{2h(\nu)}{\lambda_{1}(\nu)+\lambda_{2}(\nu)}=-q\frac{2h(\mu,\nu)}{\lambda_{1}(\nu)+\lambda_{2}(\nu)}-\tau^{\prime}(q)=q\tau^{\prime}(q)-\tau(q).$
@@
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be a dominated tuple of contractive matrices, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, and $q\neq 1$. If $\tau^{\prime}(q)$, $\tau(q)/(q-1)$, and $q\tau^{\prime}(q)-\tau(q)$ are contained in the same interval, $(0,1)$, $(1,2)$ or $(2,\infty)$, then $\tau^{\prime}(q)\leqslant\tau(q)/(q-1)$ and
@@ 5 Exact Dimensionality
for all $\mathtt{i}=i_{1}i_{2}\cdots\in\Sigma$. Here $\varphi_{\mathtt{i}}=\varphi_{i_{1}}\circ\cdots\circ\varphi_{i_{n}}$ for all $\mathtt{i}=i_{1}\cdots i_{n}\in\Sigma_{n}$ and $n\in\mathbb{N}$. It is easy to see that $\pi(\Sigma)=X$. We say that $X$ satisfies the strong separation condition if $\varphi_{i}(X)\cap\varphi_{j}(X)=\emptyset$ whenever $i\neq j$. As $\pi([\mathtt{i}])=\varphi_{\mathtt{i}}(X)$ for all $\mathtt{i}\in\Sigma_{*}$, the strong separation condition is characterized by the property that the canonical projection $\pi$ is one-to-one. We adopt the convention that whenever we speak about a self-affine set, it is automatically accompanied by the tuple of affine maps that defines it. This allows us to write, for example, “a self-affine set $X$ is dominated,” which is understood to mean that “the associated tuple $\mathsf{A}$ of matrices is dominated.”
Let $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ be a quasi-Bernoulli measure. If $X$ is dominated, then, by [6, Theorem 2.6] and [34, Theorem 2.2],
$\operatorname{dim_{loc}}(\pi_{*}\nu,\pi(\mathtt{i}))=\operatorname{\underline{dim}_{H}}(\pi_{*}\nu)\leqslant\operatorname{dim_{L}}(\nu)$ (5.1)
for $\nu$-almost all $\mathtt{i}\in\Sigma$. In other words, canonical projections of quasi-Bernoulli measures, which we call quasi-Bernoulli measures on $X$, are exact-dimensional. We say that $\mathsf{A}=(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ is irreducible if there is no line in $\mathbb{R}^{2}$ which is invariant under all of the matrices in $\mathsf{A}$. If $\mathsf{A}$ is dominated such that there is a finite set of lines in $\mathbb{R}^{2}$ invariant under all of the matrices in $\mathsf{A}$, then, by [5, Lemma 2.10], $\mathsf{A}$ is not irreducible. Therefore, if $X$ is dominated and irreducible, and satisfies the strong separation condition, then, by [2, Theorem 1.2 and the associated footnote],
@@
for all quasi-Bernoulli measures $\nu\in\mathcal{M}_{\sigma}(\Sigma)$. As observed in [5, Theorem 2.18], (5.2) holds also when $X$ is dominated and $X_{F}$ is not a singleton.
In this section, by generalizing the approach used in [16, §4.1], we study the local dimension of a quasi-Bernoulli measure on a self-affine set with respect to another measure. We first recall the following lemma.
@@
The proof is the same as that of [16, Lemma 4.2] where the result is formulated in the case of positive matrices. ∎
@@
Let $X$ be a dominated planar self-affine set satisfying the strong separation condition. If $\mu\in\mathcal{M}(\Sigma)$ is fully supported quasi-Bernoulli and $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ is ergodic and quasi-Bernoulli, and assume that
@@
satisfies $\nu\times\nu_{F}(\Gamma_{\kappa})>1-\kappa$. Since $\nu$ is quasi-Bernoulli, there exists a $T$-invariant ergodic measure $\nu^{\ast}$ on $\Sigma\times\mathbb{RP}^{1}$ which is equivalent to $\nu\times\nu_{F}$. Therefore, $\nu^{\ast}(\Gamma_{\kappa})>1-\kappa$ and Birkhoff’s ergodic theorem implies that
@@
$=\nu^{\ast}(\Gamma_{\kappa})>1-\kappa$
for $\nu^{\ast}$-almost all $(\mathtt{i},V)\in\Sigma\times\mathbb{RP}^{1}$ and hence for $\nu\times\nu_{F}$-almost all $(\mathtt{i},V)\in\Sigma\times\mathbb{RP}^{1}$. Write $r_{n}(\mathtt{i})=\alpha_{2}(\mathtt{i}|_{n})/\alpha_{1}(\mathtt{i}|_{n})$ for all $\mathtt{i}\in\Sigma$ and $n\in\mathbb{N}$, and note that, by domination, there exists $n_{0}\in\mathbb{N}$ such that $r_{n}(\mathtt{i})<r_{0}$ for all $\mathtt{i}\in\Sigma$ and $n\geqslant n_{0}$. By letting $\kappa\downarrow 0$, it follows from the above estimate that
@@
Let $\varrho_{2}>0$ be as in Lemma 5.1 and notice that, since $\alpha_{2}$ is almost-multiplicative by domination, there exists a constant $C\geqslant 1$ such that
$1$ $\geqslant\frac{\log\varrho_{2}\alpha_{2}(\mathtt{i}|_{n_{k+1}})}{\log\varrho_{2}\alpha_{2}(\mathtt{i}|_{n_{k}})}\geqslant\frac{\log\varrho_{2}\alpha_{2}(\mathtt{i}|_{n_{k}})+\log\alpha_{2}(\sigma^{n_{k}}\mathtt{i}|_{n_{k+1}-n_{k}})+\log C}{\log\varrho_{2}\alpha_{2}(\mathtt{i}|_{n_{k}})}$